# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ II.1.2 Parameterized quantum circuits (PQC)
In the NISQ era, we face significant limitations in terms of the number of qubits we can effectively use. Therefore, it’s essential to consider the quantum hardware on which our VQA runs. Constructing problem-inspired initial states can be computationally expensive when using native gates. Consequently, most ansätze developed so far fall into one of two categories: problem inspired ansätze and hardware efficient ansätze Bharti et al. (2021). The categorization depends on the structure and intended application of the ansätze. The PQC shown in FIG. 2 and FIG. 1 are examples of the second kind, hardware efficient ansätz. Note that these two circuits are also the same circuit for quantum neural network. We will now discuss both of them below.
@@
    In the problem-inspired ansätze approach, we utilize the evolution by a unitary operator, denoted as $G(t)$, with a Hermitian generator $\hat{g}~{}:$ $G(t)=e^{-i\hat{g}t}.$ These operators are derived from the properties or symmetries of the system under consideration and are used in constructing the PQC. Implementing such unitary transformations directly can be challenging. Therefore, one often employs a technique called the Suzuki-Trotter (ST) expansion or decomposition Suzuki (1976). This method provides a general approach to approximate complex and challenging-to-implement unitary transformations. In practice, the more we know about the Hamiltonian to be Trotterized, the less number of gates is needed to implement this method Kivlichan et al. (2018). Two mostly used ansätze of this kind is the unitary coupled cluster (UCC) for quantum chemistry and the quantum approximate optimization algorithm (QAOA).
@@
      Historically, the Unitary Coupled Cluster (UCC) ansätze were among the earliest proposed and implemented approaches, as discussed in reference Taube and Bartlett (2006). The core concept behind UCC is to introduce quantum correlations into the Hartree-Fock approximation. Interestingly, in 2014, it was demonstrated that representing UCC on a classical computer is computationally inefficient Yung et al. (2014). However, during the same year, researchers took advantage of quantum resources and successfully realized the UCC ansätze as a PQC using a photonic processor Peruzzo et al. (2014); Lee et al. (2018); Grimsley et al. (2019); Bharti et al. (2021).
@@
      The Quantum Approximate Optimization Algorithm (QAOA) shares strong connections with quantum annealing and adiabatic quantum computing. Adiabatic quantum computing, in brief, leverages the adiabatic theorem in quantum mechanics, which asserts that a quantum state undergoing slow variations in a time-dependent Hamiltonian will evolve only by a phase. Consequently, one can devise a Hamiltonian that evolves over time, transitioning from an easily prepared system to a system capable of encoding problems to be solved. The energy eigenstate naturally evolves from the initially prepared state to one that solves the target problem. QAOA can be viewed as an approach that resembles Trotterized adiabatic quantum computing Bharti et al. (2021); Farhi et al. (2014).
      For combinatorial optimization problems featuring strict constraints, you can introduce penalty terms in the cost function. However, this approach may not be the most efficient in practice since it can still lead to solutions that violate certain hard constraints. A variant of QAOA aimed at addressing these constraints is also discussed in previous works Farhi et al. (2014); Bharti et al. (2021).
@@
    While computational studies have demonstrated that problem inspired ansätze can yield rapid convergence to a satisfactory solution state, their experimental realization can present challenges. Quantum computing devices are subject to various experimental limitations, including factors like coherence time, limited gate fidelity, and constrained gate set. To give an example on how bad the situation is, Moll et al. (2018) illustrates that current quantum hardware is not well-suited for implementing deep and highly connected circuits, which are essential for approaches like the UCC and similar methods designed for applications beyond basic demonstrations, such as simulating more complex molecules like Hydrogen molecule. Human computation capabilities have allowed us to understand the properties of the hydrogen atom and, using manual calculations and the Born-Oppenheimer approximation, we can even compute some characteristics of hydrogen molecules. So, with problem-inspired ansätze on current devices, computers haven’t significantly expanded our understanding beyond what we already knew through traditional human-driven calculations.
    Alternatively, a class of hardware-efficient ansätze has been proposed to accommodate device constraints Kandala et al. (2017). The idea is to make good use of a limited number of quantum gates and a particular qubit connection topology. Typically, the gate set used in quantum circuits includes a two-qubit entangling gate and up to three single-qubit gates. The circuit is then built by assembling blocks of single-qubit gates and entangling gates, which are applied in parallel to multiple or all qubits. Each of these blocks is often referred to as a “layer”, and the ansatz circuit typically comprises multiple such layers.
@@
    $U=\prod_{\ell=1}^{L}W_{\ell}U_{\ell}(\theta_{\ell})=\prod_{\ell=1}^{L}W_{\ell}\exp\big{(}i\theta_{\ell}X_{\ell}\big{)}~{}.$ (2)
    The $\theta_{\ell}$ represent the variational parameters, while $U_{\ell}$ are unitary operators with $X_{\ell}$ denoting the Hermitian operators. Usually, $X_{\ell}$ takes the form of a Pauli operator, and $U_{\ell}$ consists of rotation operators that act on individual qubits. On the other hand, $W_{\ell}$ represents the multi-qubit gates responsible for entangling qubits in the circuit. For example, $W_{\ell}$ might be a CNOT gate, CZ gate, or similar multi-qubit gates commonly used in superconducting qubits Krantz et al. (2019). In line with this methodology, the “Alternating Layered Ansätze” is a specific instance of hardware-efficient ansätze. It comprises layers of single-qubit rotations and blocks of entangling gates that entangle only a local set of qubits. These entangling gate blocks are shifted every alternating layer in the circuit. The selection of these gates, their interconnections, and their order in the circuit significantly impact the portion of the Hilbert space that the ansatz explores and the rate at which it converges for a given problem Bharti et al. (2021).
@@
II.1.3 Measurements
Now, we move into the third stage of VQA, which is the measurement stage. The output of quantum circuits is a quantum state, and in order to optimize it on a classical computer, we must extract classical information from it. This involves estimating the expectation value of the objective function $\braket{\hat{\mathcal{O}}}_{U_{\theta}}$. The most direct approach is to apply a unitary operator to the quantum state, transforming it into the diagonal basis of the observable $\hat{O}$, and then calculate the probability of measuring the state with a corresponding eigenvalue of $\hat{O}$. For experimental details of this approach, see Krantz et al. (2019); HAFFNER et al. (2008). However, this direct approach is limited by the NISQ devices currently available, as transforming the state into the diagonal basis can be computationally expensive. Below, we introduce alternative methods that is more NISQ-friendly.
@@ II.1.3 Measurements
    The Pauli string: $\hat{P}=\prod_{k\in K}\bm{\sigma}(k)$ where $k$ refers to $k^{th}$ qubit, expectation value is: $\braket{\hat{P}}=\left\langle\prod_{k\in K}\sigma_{z}(k)\right\rangle_{\tilde{U}}~{},$ where $\tilde{U}$ refers to the product of rotations according to the specific form of the Pauli string. In practice, it’s only feasible to take a finite number of single-shot measurements, denoted as $N_{s}$, of the quantum state. This enables the estimation of expectation values within a finite margin of error. For more information on error of the measurements, see Huang et al. (2019).
@@
    However, this method does require ample quantum resources to implement the control unitary. An alternative approach, as proposed by the authors of Mitarai and Fujii (2019), involves decomposing $U$ into a sum of Pauli strings and making measurements with respect to these Pauli strings individually. Another approach is applicable when $U$ can be expressed as a product of unitaries, denoted as $U_{k}$, which act locally on a few qubits. This involves finding classical means to express $U_{k}$ as $U_{k}=V_{k}^{\dagger}DV_{k}$, where $D$ is a diagonal matrix and $V_{k}$ represents unitaries that can be applied to |$\ket{\psi}$ to determine the overlap. This last approach will lead us naturally to the last part of this section.
@@
    This technique leverages both the quantum and classical properties of particles. As multiple copies of quantum states’ dynamics converge into classical behaviors over time. When dealing with an unknown quantum state represented as $\rho$, the objective is to predict $M$ expectation values, denoted as $\operatorname{Tr}(\hat{O}_{i}\rho)$, where $1\leq i\leq M$ Huang et al. (2020a).
@@
    Through post-processing of the measurement results, it is possible to obtain a “classical shadow,” which serves as a classical representation of the quantum state. Research suggests that a classical shadow of size on the order of $\log{M}$ is sufficient to predict all $M$ expectation values simultaneously. This method draws inspiration from the study of shadow tomography Aaronson (2020), which will be explored further in section IV.1. As mentioned in the introduction, this is an example of the application of other topics in quantum machine learning (shadow tomography) on the narrow sensed quantum machine learning described before.
II.1.4 Optimization
Finally, we arrive at the part where we optimize the classical parameters. This optimization process does not differ significantly from other multi-variable optimization problems, so we can employ classical optimization methods Bharti et al. (2021). The challenge here, as mentioned earlier, is that in the NISQ era, we are constrained by the limited coherence time of quantum computers. This means that implementing deep analytical gradient descent circuits may not be feasible.
@@ II.1.4 Optimization
    $\mathcal{L}_{\text{MSE}}\equiv\frac{1}{2}\left(z(x_{\delta}~{};\bm{\theta})-y_{\delta}\right)^{2}~{}.$ (3)
    In machine learning, the term $z(x_{\delta}~{};\bm{\theta})$ typically represents the final output, and $y_{\delta}$ corresponds to the supervised learning data. In the context of VQA, consider a simple example where $z(x_{\delta}~{};\bm{\theta})=\braket{\psi}_{U(\theta)}$ and $y_{\delta}=0$. Therefore, minimizing the loss function is equivalent to minimizing the objective function. All the parameters are encapsulated in the vector $\bm{\theta}$.
@@
    $\theta_{i}(t+1)\equiv\theta_{i}(t)-\eta~{}\frac{\partial\mathcal{L}(\bm{\theta})}{\partial\theta_{i}}\bigg{|}_{\theta_{i}=\theta_{i}(t)}~{}.$ (4)
    This update will necessarily decrease the loss function Roberts et al. (2022). So, the iterative application of these updates will gradually bring us closer to a local minimum. However, a significant challenge with this method is the need for a substantial number of measurements. To tackle this problem, stochastic gradient descent (SGD) offers a modification to the standard parameter update rule, where it utilizes $\mathcal{L}_{\mathcal{S}_{t}}$ which serves as an approximation of the loss function. For instance, you can compute the gradient using a finite number of measurements, as described in the work by Harrow and Napp (2021). In machine learning, $\mathcal{L}_{\mathcal{S}_{t}}$ can be thought of as the loss function computed based on a subset of the training data. These subsets within the training data are randomly divided into equal-sized partitions Roberts et al. (2022).
@@
      Evolutionary Algorithms: This technique involves data sampling to estimate the gradient of expected fitness, which is then used to adjust parameters in the direction of steepest improvement Wierstra et al. (2011); Anand et al. (2021); Zhao et al. (2020).
@@
      Reinforcement Learning: In this method, reinforcement learning is initially applied to optimize parameters, as seen in the case of QAOA parameters. It employs a reinforcement learning framework to discover the most effective “policy” that maximizes the expected total discounted reward Garcia-Saez and Riu (2019); Yao et al. (2020); Wauters et al. (2020).
@@
      Sequential Minimal Optimization (SMO): In the field of machine learning, SMO has been highly effective for optimizing the parameter landscape of high-dimensional support vector machines. SMO divides the optimization task into smaller components for which analytical solutions can be derived Platt (1998); Nakanishi et al. (2020); Ostaszewski et al. (2021).
II.2 Quantum neural tangent kernel
It is important to highlight that the circuits used in quantum machine learning are identical to those employed in variational quantum algorithms (see Fig. 2). Therefore, in this section, we introduce a theory regarding the quantum neural tangent kernel (QNTK), which investigates the dynamics of VQA circuits from first principles Liu et al. (2022c). This can also be understood as an extension of the first-principle theory for classical machine learning Roberts et al. (2022).
Consider an initially prepared state, denoted as |$\ket{\Psi_{0}}$, and apply the unitary operator of the circuit to it (see Eqn. (2)): |$\ket{\phi(\bm{\theta})}=U(\bm{\theta})\ket{\Psi_{0}}=\prod_{\ell=1}^{L}W_{\ell}\exp\big{(}i\theta_{\ell}X_{\ell}\big{)}\ket{\Psi_{0}}~{}.$ Here, $\ell$ represents the layer of the quantum machine learning system. The output of the quantum neural network is then defined as: $z_{i;\delta}\equiv z_{i}(\bm{\theta},\bm{x}_{\delta})=\bra{\Psi_{0}(\bm{x}_{\delta})}U^{\dagger}\mathcal{O}_{i}U\ket{\Psi_{0}(\bm{x}_{\delta})}~{}.$ We use the mean squared loss (see Eqn.(3). We defined the error factor or residual training error $\varepsilon$ as the part in the parenthesis. The value of $y_{i;\alpha}$ is given by the supervised learning database.

### v2

@@ 2.1.2 Parameterized quantum circuits (PQC)
In the NISQ era, we face significant limitations in terms of the number of qubits we can effectively use. Therefore, it’s essential to consider the quantum hardware on which our VQA runs. Constructing problem-inspired initial states can be computationally expensive when using native gates. Consequently, most ansätze developed so far fall into one of two categories: problem inspired ansätze and hardware efficient ansätze [17]. The categorization depends on the structure and intended application of the ansätze. The PQC shown in FIG. 2 and FIG. 1 are examples of the second kind, hardware efficient ansätz. Note that these two circuits are also the same circuit for quantum neural network. We will now discuss both of them below.
@@
    In the problem-inspired ansätze approach, we utilize the evolution by a unitary operator, denoted as $G(t)$, with a Hermitian generator $\hat{g}\leavevmode\nobreak\ :$ $G(t)=e^{-i\hat{g}t}.$ These operators are derived from the properties or symmetries of the system under consideration and are used in constructing the PQC. Implementing such unitary transformations directly can be challenging. Therefore, one often employs a technique called the Suzuki-Trotter (ST) expansion or decomposition [120]. This method provides a general approach to approximate complex and challenging-to-implement unitary transformations. In practice, the more we know about the Hamiltonian to be Trotterized, the less number of gates is needed to implement this method [74]. Two mostly used ansätze of this kind is the unitary coupled cluster (UCC) for quantum chemistry and the quantum approximate optimization algorithm (QAOA).
@@
      Historically, the Unitary Coupled Cluster (UCC) ansätze were among the earliest proposed and implemented approaches, as discussed in reference [123]. The core concept behind UCC is to introduce quantum correlations into the Hartree-Fock approximation. Interestingly, in 2014, it was demonstrated that representing UCC on a classical computer is computationally inefficient [137]. However, during the same year, researchers took advantage of quantum resources and successfully realized the UCC ansätze as a PQC using a photonic processor [108, 82, 49, 17].
@@
      The Quantum Approximate Optimization Algorithm (QAOA) shares strong connections with quantum annealing and adiabatic quantum computing. Adiabatic quantum computing, in brief, leverages the adiabatic theorem in quantum mechanics, which asserts that a quantum state undergoing slow variations in a time-dependent Hamiltonian will evolve only by a phase. Consequently, one can devise a Hamiltonian that evolves over time, transitioning from an easily prepared system to a system capable of encoding problems to be solved. The energy eigenstate naturally evolves from the initially prepared state to one that solves the target problem. QAOA can be viewed as an approach that resembles Trotterized adiabatic quantum computing [17, 44].
      For combinatorial optimization problems featuring strict constraints, you can introduce penalty terms in the cost function. However, this approach may not be the most efficient in practice since it can still lead to solutions that violate certain hard constraints. A variant of QAOA aimed at addressing these constraints is also discussed in previous works [44, 17].
@@
    While computational studies have demonstrated that problem inspired ansätze can yield rapid convergence to a satisfactory solution state, their experimental realization can present challenges. Quantum computing devices are subject to various experimental limitations, including factors like coherence time, limited gate fidelity, and constrained gate set. To give an example on how bad the situation is, [100] illustrates that current quantum hardware is not well-suited for implementing deep and highly connected circuits, which are essential for approaches like the UCC and similar methods designed for applications beyond basic demonstrations, such as simulating more complex molecules like Hydrogen molecule. Human computation capabilities have allowed us to understand the properties of the hydrogen atom and, using manual calculations and the Born-Oppenheimer approximation, we can even compute some characteristics of hydrogen molecules. So, with problem-inspired ansätze on current devices, computers haven’t significantly expanded our understanding beyond what we already knew through traditional human-driven calculations.
    Alternatively, a class of hardware-efficient ansätze has been proposed to accommodate device constraints [71]. The idea is to make good use of a limited number of quantum gates and a particular qubit connection topology. Typically, the gate set used in quantum circuits includes a two-qubit entangling gate and up to three single-qubit gates. The circuit is then built by assembling blocks of single-qubit gates and entangling gates, which are applied in parallel to multiple or all qubits. Each of these blocks is often referred to as a “layer”, and the ansatz circuit typically comprises multiple such layers.
@@
    $U=\prod_{\ell=1}^{L}W_{\ell}U_{\ell}(\theta_{\ell})=\prod_{\ell=1}^{L}W_{\ell}\exp\big{(}i\theta_{\ell}X_{\ell}\big{)}\leavevmode\nobreak\ .$ (2)
    The $\theta_{\ell}$ represent the variational parameters, while $U_{\ell}$ are unitary operators with $X_{\ell}$ denoting the Hermitian operators. Usually, $X_{\ell}$ takes the form of a Pauli operator, and $U_{\ell}$ consists of rotation operators that act on individual qubits. On the other hand, $W_{\ell}$ represents the multi-qubit gates responsible for entangling qubits in the circuit. For example, $W_{\ell}$ might be a CNOT gate, CZ gate, or similar multi-qubit gates commonly used in superconducting qubits [77]. In line with this methodology, the “Alternating Layered Ansätze” is a specific instance of hardware-efficient ansätze. It comprises layers of single-qubit rotations and blocks of entangling gates that entangle only a local set of qubits. These entangling gate blocks are shifted every alternating layer in the circuit. The selection of these gates, their interconnections, and their order in the circuit significantly impact the portion of the Hilbert space that the ansatz explores and the rate at which it converges for a given problem [17].
@@
2.1.3 Measurements
Now, we move into the third stage of VQA, which is the measurement stage. The output of quantum circuits is a quantum state, and in order to optimize it on a classical computer, we must extract classical information from it. This involves estimating the expectation value of the objective function $\braket{\hat{\mathcal{O}}}_{U_{\theta}}$. The most direct approach is to apply a unitary operator to the quantum state, transforming it into the diagonal basis of the observable $\hat{O}$, and then calculate the probability of measuring the state with a corresponding eigenvalue of $\hat{O}$. For experimental details of this approach, see [77, 52]. However, this direct approach is limited by the NISQ devices currently available, as transforming the state into the diagonal basis can be computationally expensive. Below, we introduce alternative methods that is more NISQ-friendly.
@@ 2.1.3 Measurements
    The Pauli string: $\hat{P}=\prod_{k\in K}\bm{\sigma}(k)$ where $k$ refers to $k^{th}$ qubit, expectation value is: $\braket{\hat{P}}=\left\langle\prod_{k\in K}\sigma_{z}(k)\right\rangle_{\tilde{U}}\leavevmode\nobreak\ ,$ where $\tilde{U}$ refers to the product of rotations according to the specific form of the Pauli string. In practice, it’s only feasible to take a finite number of single-shot measurements, denoted as $N_{s}$, of the quantum state. This enables the estimation of expectation values within a finite margin of error. For more information on error of the measurements, see [61].
@@
    However, this method does require ample quantum resources to implement the control unitary. An alternative approach, as proposed by the authors of [99], involves decomposing $U$ into a sum of Pauli strings and making measurements with respect to these Pauli strings individually. Another approach is applicable when $U$ can be expressed as a product of unitaries, denoted as $U_{k}$, which act locally on a few qubits. This involves finding classical means to express $U_{k}$ as $U_{k}=V_{k}^{\dagger}DV_{k}$, where $D$ is a diagonal matrix and $V_{k}$ represents unitaries that can be applied to |$\ket{\psi}$ to determine the overlap. This last approach will lead us naturally to the last part of this section.
@@
    This technique leverages both the quantum and classical properties of particles. As multiple copies of quantum states’ dynamics converge into classical behaviors over time. When dealing with an unknown quantum state represented as $\rho$, the objective is to predict $M$ expectation values, denoted as $\operatorname{Tr}(\hat{O}_{i}\rho)$, where $1\leq i\leq M$ [62].
@@
    Through post-processing of the measurement results, it is possible to obtain a “classical shadow,” which serves as a classical representation of the quantum state. Research suggests that a classical shadow of size on the order of $\log{M}$ is sufficient to predict all $M$ expectation values simultaneously. This method draws inspiration from the study of shadow tomography [3], which will be explored further in section 4.1. As mentioned in the introduction, this is an example of the application of other topics in quantum machine learning (shadow tomography) on the narrow sensed quantum machine learning described before.
2.1.4 Optimization
Finally, we arrive at the part where we optimize the classical parameters. This optimization process does not differ significantly from other multi-variable optimization problems, so we can employ classical optimization methods [17]. The challenge here, as mentioned earlier, is that in the NISQ era, we are constrained by the limited coherence time of quantum computers. This means that implementing deep analytical gradient descent circuits may not be feasible.
@@ 2.1.4 Optimization
    $\mathcal{L}_{\text{MSE}}\equiv\frac{1}{2}\left(z(x_{\delta}\leavevmode\nobreak\ ;\bm{\theta})-y_{\delta}\right)^{2}\leavevmode\nobreak\ .$ (3)
    In machine learning, the term $z(x_{\delta}\leavevmode\nobreak\ ;\bm{\theta})$ typically represents the final output, and $y_{\delta}$ corresponds to the supervised learning data. In the context of VQA, consider a simple example where $z(x_{\delta}\leavevmode\nobreak\ ;\bm{\theta})=\braket{\psi}_{U(\theta)}$ and $y_{\delta}=0$. Therefore, minimizing the loss function is equivalent to minimizing the objective function. All the parameters are encapsulated in the vector $\bm{\theta}$.
@@
    $\theta_{i}(t+1)\equiv\theta_{i}(t)-\eta\leavevmode\nobreak\ \frac{\partial\mathcal{L}(\bm{\theta})}{\partial\theta_{i}}\bigg{|}_{\theta_{i}=\theta_{i}(t)}\leavevmode\nobreak\ .$ (4)
    This update will necessarily decrease the loss function [112]. So, the iterative application of these updates will gradually bring us closer to a local minimum. However, a significant challenge with this method is the need for a substantial number of measurements. To tackle this problem, stochastic gradient descent (SGD) offers a modification to the standard parameter update rule, where it utilizes $\mathcal{L}_{\mathcal{S}_{t}}$ which serves as an approximation of the loss function. For instance, you can compute the gradient using a finite number of measurements, as described in the work by [57]. In machine learning, $\mathcal{L}_{\mathcal{S}_{t}}$ can be thought of as the loss function computed based on a subset of the training data. These subsets within the training data are randomly divided into equal-sized partitions [112].
@@
      Evolutionary Algorithms: This technique involves data sampling to estimate the gradient of expected fitness, which is then used to adjust parameters in the direction of steepest improvement [134, 6, 139].
@@
      Reinforcement Learning: In this method, reinforcement learning is initially applied to optimize parameters, as seen in the case of QAOA parameters. It employs a reinforcement learning framework to discover the most effective “policy” that maximizes the expected total discounted reward [46, 136, 130].
@@
      Sequential Minimal Optimization (SMO): In the field of machine learning, SMO has been highly effective for optimizing the parameter landscape of high-dimensional support vector machines. SMO divides the optimization task into smaller components for which analytical solutions can be derived [109, 101, 105].
Also, in [16], the authors present variational quantum algorithms for real and imaginary time evolution that are more efficient than previous proposals, where they use Sequential Minimal Optimization for time evolution.
2.2 Quantum neural tangent kernel
It is important to highlight that the circuits used in quantum machine learning are identical to those employed in variational quantum algorithms (see Fig. 2). Therefore, in this section, we introduce a theory regarding the quantum neural tangent kernel (QNTK), which investigates the dynamics of VQA circuits from first principles [87]. This can also be understood as an extension of the first-principle theory for classical machine learning [112].
Consider an initially prepared state, denoted as |$\ket{\Psi_{0}}$, and apply the unitary operator of the circuit to it (see Eqn. (2)): |$\ket{\phi(\bm{\theta})}=U(\bm{\theta})\ket{\Psi_{0}}=\prod_{\ell=1}^{L}W_{\ell}\exp\big{(}i\theta_{\ell}X_{\ell}\big{)}\ket{\Psi_{0}}\leavevmode\nobreak\ .$ Here, $\ell$ represents the layer of the quantum machine learning system. The output of the quantum neural network is then defined as: $z_{i;\delta}\equiv z_{i}(\bm{\theta},\bm{x}_{\delta})=\bra{\Psi_{0}(\bm{x}_{\delta})}U^{\dagger}\mathcal{O}_{i}U\ket{\Psi_{0}(\bm{x}_{\delta})}\leavevmode\nobreak\ .$ We use the mean squared loss (see Eqn.(3). We defined the error factor or residual training error $\varepsilon$ as the part in the parenthesis. The value of $y_{i;\alpha}$ is given by the supervised learning database.